OpenStudy (anonymous):

-4ln2x=-26 solve each equation or inequality. round to the nearest ten thousandth.

OpenStudy (mertsj):

\[\ln 2x=\frac{13}{2}\]

OpenStudy (mertsj):

Now put it in exponential form

OpenStudy (anonymous):

Wait, I'm confused because so far I have \[\ln2x ^{-4}=-26\]

OpenStudy (mertsj):

Ok. It would be much easier to just divide both sides by -4 but if you must...then \[\ln (2x)^{-4}=-26\]

OpenStudy (mertsj):

\[\ln \frac{1}{(2x)^4}=-26\]

OpenStudy (mertsj):

Now put it in exponential form.

OpenStudy (anonymous):

I need an answer put to the nearest ten thousandth

OpenStudy (mertsj):

Then put it in exponential form.

OpenStudy (mertsj):

What is the base?

OpenStudy (anonymous):

\[e ^{13}=x ^{-4}\]

OpenStudy (mertsj):

\[e ^{-26}=\frac{1}{(2x)^4}\]

OpenStudy (mertsj):

So \[(2x)^4=\frac{1}{e ^{-26}}=e ^{26}\]

OpenStudy (mertsj):

Take the 4th root of both sides and divide the answer by 2

OpenStudy (mertsj):